# Path Eccentricity and Forbidden Induced Subgraphs

### v1

@@ 2 Background on Structural Domination
Suppose now that $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. If $H$ admits a vertex $u$ of degree $3$, then, since $H$ is an induced subgraph of $T_{k}$, two neighbors of $u$ are adjacent, forming a cycle. However, this is impossible, since $H$ is also a subgraph of the acyclic graph $S_{k}$. Thus $H$ has maximum degree at most $2$. This implies that $H$ is an induced subgraph of either the graph obtained by removing the vertex of degree $3$ from $S_{k}$, which is isomorphic to $3P_{k}$; or the graph obtained by removing a neighbor of the vertex of degree $3$ from $S_{k}$, which is isomorphic to $P_{2k+1}+P_{k-1}$. ∎
@@
Suppose first that $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. By Lemma 3.1, $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. This implies that every connected $H$-free graph $G$ is $\{S_{k},T_{k}\}$-free and, hence, satisfies $\operatorname{\mathbf{pe}}(G)<k$ thanks to Theorem 1.4.
Suppose now that $H$ is not an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. Then, again by Lemma 3.1, $H$ is not an induced subgraph of either $S_{k}$ or $T_{k}$. Let $R_{k}$ be a graph in $\{S_{k},T_{k}\}$ that does not admit $H$ as induced subgraph. Then $R_{k}$ is a connected $H$-free graph and ${\operatorname{\mathbf{pe}}(R_{k})\geqslant k}$ (by Theorem 1.4). ∎
4 Implications for the *-C1P Property (Theorem 1.6)
@@
Since the argument are the same for both graphs, we use the same notations for their vertices. Let $a^{\prime}$, $b^{\prime}$ and $c^{\prime}$ be the three leaves in $S_{2}$ (respectively $T_{1}$), $a$, $b$ and $c$ their respective neighbors and let $x$ the central vertex of $S_{2}$ (see Figure 5). Suppose that $S_{2}$ (respectively $T_{1}$) has the *-C1P. Then there exists an ordering $\sigma$ on the vertices such that for each vertex $u$, there exists $N_{u}\in\{N[u],N(u)\}$ (in particular, $N(u)\subseteq N_{u}\subseteq N[u]$), such that $N_{u}$ consists of consecutive vertices in the ordering $\sigma$. Observe that in both graphs, the sets $N_{a}$, $N_{b}$, and $N_{c}$ are pairwise intersecting. Since these sets correspond to consecutive sets in the ordering $\sigma$, one of those sets, w.l.o.g., $N_{a}$, is contained in the union of the others, here $N_{b}$ and $N_{c}$. However, $a^{\prime}$ belongs to $N(a)\subseteq N_{a}$ and does not belong to $N[b]\cup N[c]$, and, hence, does also not belong to $N_{b}\cup N_{c}$, a contradiction. ∎
Figure 5: Notation for the vertices of $S_{2}$ and $T_{1}$ in Lemma 4.1, with the open neighborhoods of $a$, $b$, and $c$.
@@
It follows directly from the definition of *-C1P that if $G$ is a graph that has the *-C1P and $H$ is an induced subgraph of $G$, then $H$ also has the *-C1P. Thus, by Lemma 4.1, graphs with *-C1P are $\{S_{2},T_{1}\}$-free. Therefore, they are also $\{S_{2},T_{2}\}$-free and, hence, have path eccentricity at most 1 by Theorem 1.4. ∎
5 Algorithmic Proof (Theorem 1.8)
In this section, we use the following notation regarding paths in graphs. Given a path $P$ of extremities $u,v$ and vertices $x,y$ adjacent respectively to $u,v$, we denote by $xPy$ the path obtained by adding the edges $xu$, $vy$ to $P$. Similarly, we denote by $xP$ and $Py$ the paths obtained by adding respectively only $xu$ or only $vy$ to $P$.
@@
Observe that neither $u$ nor $v$ has a neighbor outside of $P$, as $P$ is a longest path. Hence, $w$ is an internal vertex of $P$, implying in particular that $u$, $v$, and $w$ are three distinct vertices of $P$. Denote by $u^{\prime}$, $v^{\prime}$ the neighbor of $u$, resp. $v$ on $P$. Observe that $u^{\prime}\neq v$ and $v^{\prime}\neq u$. Let $P^{*}$ be the subgraph of $P$ obtained from deleting $u$ and $v$ from $P$. In case $y$ is adjacent to any of $u^{\prime},v^{\prime}$, we define $P^{\prime}$ as $xyP^{*}v$ or $uP^{*}yx$. In both cases, $P^{\prime}$ has a greater length than $P$, a contradiction. Therefore, $x$ and $y$ are non-adjacent to any of $u$, $v$, $u^{\prime}$, and $v^{\prime}$. Moreover, as $y$ is adjacent to $P$, the vertices $u^{\prime}$, $w$, and $v^{\prime}$ are pairwise distinct, implying that $|V(P)|\geqslant 5$.
@@
The above argument shows that we can modify a given longest path $P$ at most $|V(G)|$ times until we obtain a longest path that is dominating.
It remains to discuss the running time in the case we start the process above with some arbitrary (not necessarily longest) path. Then, in each step of the argument above, we modify the current path $P$ to a path $P^{\prime}$ that is either longer than $P$ or has the same length but dominates more vertices. Therefore, we run the standard BFS algorithm checking whether path $P$ is already dominating. If not, it finds a vertex $x$ at distance $2$ from $P$ together with a connecting path back to $P$ and we proceed with the algorithm finding $P^{\prime}$ and repeat. BFS runs in time $\mathcal{O}(n+m)$. Our proof guarantees at most $n^{2}$ steps, as in each step we either increase the size of the dominated set by at least $1$, or we do not decrease it but we increase the length of path $P$. ∎
@@ 6 Concluding Remarks and Open Questions
Our main result, Theorem 1.4, shows that for every integer $k\geqslant 1$, each connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. This raises the question regarding the complexity of computing the path eccentricity of such a graph. While the problem is clearly polynomial for $k=1$, it turns out that the problem is NP-hard for all $k\geqslant 2$. This follows from the fact that the problem of determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96], which are known to be $2P_{2}$-free and, hence, $\{S_{2},T_{2}\}$-free. This, however, says nothing about the complexity of computing a path witnessing the upper bound on path eccentricity guaranteed by Theorem 1.4.
@@
Is there an algorithmic version of Theorem 1.4? More precisely: for a fixed integer $k\geqslant 1$ and a connected $\{S_{k},T_{k}\}$-free graph $G$, can we find in polynomial time a path in $G$ with eccentricity less than $k$?
While 6.1 is known to have an affirmative answer for $k=1$ (see [She91, BDK00]), it is, to the best of our knowledge, open for all $k\geqslant 2$. In fact, even a weaker version of 6.1, finding a path of eccentricity less than $k$ in cases covered by Corollary 1.7, is open, except for the case of $3P_{2}$-free graphs, which is covered by Theorem 1.8.
Another result of this paper (Theorem 1.8) is that every connected $3P_{2}$-free graph $G$ has a longest path that is dominating. This result, along with the observation that there exist graphs that are not $3P_{2}$-free such that every connected induced subgraph has a longest path that is dominating (consider, for instance, paths and cycles), raise the question about a characterization of such graphs.
@@
Is it true that for every integer $k\geqslant 2$, every connected $\{S_{k},T_{k}\}$-free graph has a minimum connected distance-$(k-1)$ dominating set inducing a $\{S_{1},T_{1}\}$-free graph?

### v2

@@ 2 Background on Structural Domination
Suppose now that $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. If $H$ admits a vertex $u$ of degree $3$, then, since $H$ is an induced subgraph of $T_{k}$, two neighbors of $u$ are adjacent, forming a cycle. However, this is impossible, since $H$ is also a subgraph of the acyclic graph $S_{k}$. Thus $H$ has maximum degree at most $2$. This implies that $H$ is an induced subgraph of either the graph obtained by removing the vertex of degree $3$ from $S_{k}$, which is isomorphic to $3P_{k}$, or the graph obtained by removing a neighbor of the vertex of degree $3$ from $S_{k}$, which is isomorphic to $P_{2k+1}+P_{k-1}$. ∎
@@
Suppose first that $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. By Lemma˜3.1, $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. This implies that every connected $H$-free graph $G$ is $\{S_{k},T_{k}\}$-free and, hence, satisfies $\operatorname{\mathbf{pe}}(G)<k$ thanks to Theorem˜1.4.
Suppose now that $H$ is not an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. Then, again by Lemma˜3.1, $H$ is not an induced subgraph of either $S_{k}$ or $T_{k}$. Let $R_{k}$ be a graph in $\{S_{k},T_{k}\}$ that does not admit $H$ as an induced subgraph. Then $R_{k}$ is a connected $H$-free graph and $\operatorname{\mathbf{pe}}(R_{k})\geqslant k$} (by Theorem˜1.4). ∎
4 Implications for the *-C1P Property (Theorem˜1.6)
@@
Since the argument is the same for both graphs, we use the same notations for their vertices. Let $a^{\prime}$, $b^{\prime}$, and $c^{\prime}$ be the three leaves in $S_{2}$ (respectively $T_{1}$), $a$, $b$, and $c$ their respective neighbors and let $x$ the central vertex of $S_{2}$ (see Figure˜6). Suppose that $S_{2}$ (respectively $T_{1}$) has the *-C1P. Then there exists an ordering $\sigma$ on the vertices such that for each vertex $u$, there exists $N_{u}\in\{N[u],N(u)\}$ (in particular, $N(u)\subseteq N_{u}\subseteq N[u]$), such that $N_{u}$ consists of consecutive vertices in the ordering $\sigma$. Observe that in both graphs, the sets $N_{a}$, $N_{b}$, and $N_{c}$ are pairwise intersecting. Since these sets correspond to consecutive sets in the ordering $\sigma$, one of those sets, w.l.o.g., $N_{a}$, is contained in the union of the others, here $N_{b}$ and $N_{c}$. However, $a^{\prime}$ belongs to $N(a)\subseteq N_{a}$ and does not belong to $N[b]\cup N[c]$, and, hence, also does not belong to $N_{b}\cup N_{c}$, a contradiction. ∎
Figure 6: Notation for the vertices of $S_{2}$ and $T_{1}$ in Lemma˜4.1, with the open neighborhoods of $a$, $b$, and $c$.
@@
It follows directly from the definition of *-C1P that if $G$ is a graph that has the *-C1P and $H$ is an induced subgraph of $G$, then $H$ also has the *-C1P. Thus, by Lemma˜4.1, graphs with *-C1P are $\{S_{2},T_{1}\}$-free. Therefore, they are also $\{S_{2},T_{2}\}$-free and, hence, have path eccentricity at most 1 by Theorem˜1.4. ∎
5 Algorithmic Proofs
We next prove our main result regarding $3P_{2}$-free graphs. While the statement of the theorem is not algorithmic, the proof is, in the sense that it yields an efficient algorithm transforming a given longest path in a $3P_{2}$-free graph into a longest path that is dominating. (Recall, however, that identifying a longest path in a given $3P_{2}$-free graph is an NP-hard problem.)
In this section, we use the following notation regarding paths in graphs. Given a path $P$ with extremities $u,v$ and vertices $x,y$ adjacent respectively to $u,v$, we denote by $xPy$ the path obtained by adding the edges $xu$, vyvy to $P$. Similarly, we denote by $xP$ and $Py$ the paths obtained by adding respectively only $xu$ or only $vy$ to $P$. For any integer $k\geqslant 0$ and path $P$ in a graph $G$, we denote by $N_{k}[P]$ the set of vertices at distance at most $k$ from some vertex of $P$. For $k=1$, we write $N[P]$ instead of $N_{1}[P]$. We say that a path $P$ $k$-dominates a vertex $x\in V(G)$ if $x\in N_{k}[P]$. For $k=1$, we simply say that $P$ dominates $x$.
@@ 5 Algorithmic Proofs
Observe that neither $u$ nor $v$ has a neighbor outside of $P$, as $P$ is a longest path. Hence, $w$ is an internal vertex of $P$, implying in particular that $u$, $v$, and $w$ are three distinct vertices of $P$. Denote by $u^{\prime}$, $v^{\prime}$ the neighbor of $u$, respectively $v$ on $P$. Observe that $u^{\prime}\neq v$ and $v^{\prime}\neq u$. Let $P^{*}$ be the subgraph of $P$ obtained by deleting $u$ and $v$ from $P$. In the case $y$ is adjacent to any of $u^{\prime},v^{\prime}$, we define $P^{\prime}$ as $xyP^{*}v$ or $uP^{*}yx$. In both cases, $P^{\prime}$ has a greater length than $P$, a contradiction. Therefore, $x$ and $y$ are non-adjacent to any of $u$, $v$, $u^{\prime}$, and $v^{\prime}$. Moreover, as $y$ is adjacent to $P$, the vertices $u^{\prime}$, $w$, and $v^{\prime}$ are pairwise distinct, implying that $|V(P)|\geqslant 5$.
@@
The above argument shows that we can modify a given longest path $P$ at most $|V(G)|$ times until we obtain a longest path that is dominating. ∎
Before we proceed with the proof of Theorem˜1.9, we will need the following lemma from [DDH+24] regarding a sufficient condition for the presence of an induced subgraph isomorphic to a graph from $\mathcal{S}$, $\mathcal{T}$, or $\mathcal{M}$ (recall the definitions on p. 5 and Figure˜5).
Lemma 5.1 (Dallard et al. [DDH+24, Lemma 1]).
Let $G$ be a graph and $I$ be an independent set in $G$ with $|I|=3$. If there exists a component $C$ of $G\setminus I$ such that $I\subseteq N(C)$, then there exists an induced subgraph $H$ of $G[N[C]]$ such that $H\in\mathcal{S}\cup\mathcal{T}\cup\mathcal{M}$ and $I$ is exactly the set of extremities of $H$.
Moreover, from the proof of the above lemma (see Lemma 6.1,[BDD+25]), we can deduce that $H$ can be found in $\mathcal{O}(n+m)$ time.
See 1.9
Proof.
Let $P$ be an arbitrary path in $G$ and $u,v$ be the extremities of $P$.
The algorithm proceeds recursively as follows. While $N_{k-1}[P]\neq V(G)$, we construct a path $P^{\prime}$ such that:
  • (i)
    $N_{k-1}[P]\subsetneq N_{k-1}[P^{\prime}]$ (corresponding to ˜5.3), or
  • (ii)
    $N_{k-1}[P]\subseteq N_{k-1}[P^{\prime}]$ and $|V(P^{\prime})|<|V(P)|$ (corresponding to ˜5.2).
After at most $n$ updates of type (i), $N_{k-1}[P]=V(G)$, that is, a path of with eccentricity less than $k$ in $G$ is obtained. Before an update of type (i), we create a nice setting for such update by applying updates of type (ii) as many times as possible (which is at most $n$ times). If at some point it is not possible to execute an update of type (i), we exhibit an induced subgraph isomorphic to either $S_{k}$, $T_{k}$, or a graph in $\mathcal{M}_{k}$. In what follows, we show that one update takes $\operatorname{\mathcal{O}}(n+m)$. This will imply that the running time of the algorithm is $\operatorname{\mathcal{O}}(n^{2}(n+m))$.
Claim 5.2.
If $u\neq v$, then at least one of the following holds.
  • $N_{k-1}[P-u]=N_{k-1}[P]$,
  • $N_{k-1}[P-v]=N_{k-1}[P]$, or
  • there exist two distinct vertices $u^{\prime}$ and $v^{\prime}$ in $G$ and two paths $P_{u}$ and $P_{v}$ with extremities $u,u^{\prime}$ and $v,v^{\prime}$ respectively, such that $V(P_{u})\cap V(P_{v})=\emptyset$, $V(P_{u})\cap V(P)=\{u\}$, $V(P_{v})\cap V(P)=\{v\}$, and the sets $V(P_{u})$ and $V(P_{v})$ each induce a $P_{k}$ in $G$. Moreover, the only edge between $V(P_{u})\setminus\{u\}$ (respectively $V(P_{v})\setminus\{v\}$) and $V(P)$ is the edge between $u$ (respectively $v$) and its neighbor along $P_{u}$ (respectively $P_{v}$).
Distinguishing which of these cases occurs and, in the case of the last outcome, a certificate for it can be found in time $\operatorname{\mathcal{O}}(n+m)$.
Proof of ˜5.2.
Suppose that $u\neq v$. If $N_{k-1}[P-u]\subsetneq N_{k-1}[P]$, then there exists a vertex $u^{\prime}$ at distance k1k-1 from $u$ such that $u^{\prime}$ is at distance $k$ from PuP-u. Let $P_{u}$ be a path witnessing distance k1k-1 between $u^{\prime}$ and $u$. Then $P_{u}$ induces a $P_{k}$ in $G$ and is internally disjoint from $P$ since $u^{\prime}$ is at distance $k$ from PuP-u. Moreover, the only edge between $V(P_{u})\setminus\{u\}$ and $V(P)$ is the edge between $u$ and its neighbor along $P_{u}$. If $N_{k-1}[P-v]\subsetneq N_{k-1}[P]$ also holds, we construct similarly $v^{\prime}$ and $P_{v}$ with analogous properties.
Suppose that $P_{u}$ and $P_{v}$ intersect in some vertex $x$. Without loss of generality, the distance from $u^{\prime}$ to $x$ along $P_{u}$ is at most the distance from $v^{\prime}$ to $x$ along $P_{v}$. Then, following first $P_{u}$ between $u^{\prime}$ and $x$ and then following $P_{v}$ between $x$ and $v$, we get a path from $u^{\prime}$ to $v$ at most as long as $P_{v}$, which has length k1k-1, thus $u^{\prime}$ is at distance at most k1k-1 from $v\in V(P-u)$, a contradiction.
The time complexity follows, as we first check whether one of the first two outcomes holds, comparing outcomes of the BFS algorithms starting with $P$, PuP-u, and PvP-v, respectively. If all those outcomes are different, they provide us with $u^{\prime}$ and $v^{\prime}$ and the respective paths $P_{u}$ and $P_{v}$. ∎
Assume that $u\neq v$. Note that if the first or the second case of ˜5.2 holds, then we can replace $P$ by either PuP-u or PvP-v, and get a path strictly smaller with the same neighborhood at distance k1k-1. Repeating this at most $|V(P)|<n$ times, we obtain $P^{\prime}$ such that $N_{k-1}[P^{\prime}]=N_{k-1}[P]$ and $|V(P^{\prime})|\leqslant|V(P)|$, and either $P^{\prime}$ is reduced to a single vertex, or we get the existence of $u^{\prime},v^{\prime},P_{u},P_{v}$ as defined in ˜5.2 (with $P^{\prime}$ in place of $P$). We then replace $P$ by $P^{\prime}$.
We can now assume that either $P$ is reduced to a single vertex, or there exist $u^{\prime},v^{\prime},P_{u},P_{v}$ as defined in ˜5.2. We can also assume that $N_{k-1}[P]\neq V(G)$, since otherwise $P$ is a path of eccentricity less than $k$ and we can return it. Thus, there exists a vertex $x$ at distance $k$ from $P$. Therefore, there is an induced $P_{k}$ in $G$, denoted $P_{x}$, with extremities $x$ and $y$ such that $y$ is the only vertex of $P_{x}$ adjacent to some vertex of $P$.
Claim 5.3.
Let
$S:=\begin{cases}V(P_{x})\cup V(P_{u})\cup V(P_{v}),\hfill\text{~~~ if $|V(P)|\neq 1$.}\\ V(P_{x}),\hfill\text{~~~otherwise.}\end{cases}$
Then, either $S$ induces a $3P_{k}$ or we can construct a path $P^{\prime}$ in $S\cup V(P)$ such that $N_{k-1}[P]\subsetneq N_{k-1}[P^{\prime}]$. A certificate for one of the outcomes can be found in time $\operatorname{\mathcal{O}}(n+m)$.
Proof of ˜5.3.
If $P$ is reduced to a single vertex, then, by the construction of $P_{x}$, the set $V(P)\cup V(P_{x})$ induces a path that contains $V(P)\cup\{x\}$, implying that $N_{k-1}[P]\subsetneq N_{k-1}[P^{\prime}]$, since $x\notin N_{k-1}[P]$, so the claim holds.
Suppose now that $u\neq v$ (thus $u^{\prime},v^{\prime},P_{u}$, and $P_{v}$ are defined as in ˜5.2). Recall that each path $P_{u},P_{v},P_{x}$ taken individually is an induced $P_{k}$ in $G$. If $V(P_{x})\cup V(P_{u})\cup V(P_{v})$ induces a $3P_{k}$ then return this outcome. Otherwise, this means that either they share some vertices or that there exists an edge with endpoints in two of those three paths. In each case, we will construct a path $P^{\prime}$ in $V(P_{x})\cup V(P_{u})\cup V(P_{v})\cup V(P)$ that contains $V(P)\cup\{x\}$, implying that $N_{k-1}[P]\subsetneq N_{k-1}[P^{\prime}]$, since $x\notin N_{k-1}[P]$.
Suppose first that $P_{x}$, $P_{u}$, and $P_{v}$ are not vertex-disjoint. Since by ˜5.2, $P_{u}$ and $P_{v}$ are vertex-disjoint, $V(P_{x})$ intersects $V(P_{u})\cup V(P_{v})$. Let $y^{\prime}$ be the vertex of $P_{x}$ closest to $x$ along $P_{x}$ that belongs to $V(P_{u})\cup V(P_{v})$. By symmetry, we may assume that $y^{\prime}\in V(P_{u})$, and we denote $P_{x}^{\prime}$ and $P_{u}^{\prime}$ the respective subpaths of $P_{x}$ with extremities $x,y^{\prime}$ and of $P_{u}$ with extremities $y^{\prime},u$. Then we construct our wanted $P^{\prime}$ as the concatenation of $P_{x}^{\prime}$, $P_{u}^{\prime}$, and $P$. We can thus assume that $V(P_{x})$, $V(P_{u})$, and $V(P_{v})$ are pairwise disjoint.
If there is an edge $e$ between a vertex $y^{\prime}\in V(P_{x})$ and a vertex $u^{\prime\prime}\in V(P_{u})$, we define $P_{x}^{\prime}$ and $P_{u}^{\prime}$ as the respective subpaths of $P_{x}$ with extremities $x,y^{\prime}$ and of $P_{u}$ with extremities $u^{\prime\prime},u$. Then we construct our wanted $P^{\prime}$ as the concatenation of $P_{x}^{\prime}$, $e$, $P_{u}^{\prime}$, and $P$. The case where $e$ is between a vertex of $P_{x}$ and $P_{v}$ is obtained by symmetry. We can now assume that there is no edge between $V(P_{x})$ and $V(P_{u})\cup V(P_{v})$.
Finally, suppose there is an edge $e$ between a vertex $u^{\prime\prime}\in V(P_{u})$ and a vertex $v^{\prime\prime}\in V(P_{v})$, we define $P_{u}^{\prime}$ and $P_{v}^{\prime}$ as the respective subpaths of $P_{u}$ with extremities $u^{\prime\prime},u$ and of $P_{v}$ with extremities $v,v^{\prime\prime}$. Furthermore, let $w$ be a neighbor of $y$ in $P$. Since there is no edge between $V(P_{u})$ and $V(P_{x})$, we have $w\neq u$. Thus, there exists $w^{\prime}$, the neighbor of $w$ in $P$ toward $u$. Let $Q$ and $Q^{\prime}$ be two subpaths of $P$ with extremities $w,v$ and $u,w^{\prime}$ respectively; note that $V(Q)\cup V(Q^{\prime})=V(P)$. Then we construct our wanted $P^{\prime}$ as the concatenation of $P_{x}$, $yw$, $Q$, $P_{v}^{\prime}$, $e$, $P_{u}^{\prime}$, and $Q^{\prime}$.
The time complexity follows, as any edge or a shared vertex can be discovered within $S$ by scanning the closed neighborhood of $P_{x}$ and $P_{u}$, in which case we can output $P^{\prime}$. ∎
Now suppose that ˜5.3 returned the outcome that $S=V(P_{x})\cup V(P_{u})\cup V(P_{v})$ induces a $3P_{k}$. Let $G^{\prime}=G[V(P)\cup\{y\}]$. Then $I=\{u,v,y\}$ is an independent set in $G^{\prime}$, since $S$ induces a $3P_{k}$. Let $C$ be the subgraph of $G^{\prime}$ induced by $V(P)\setminus I$. Then $I\subset N(C)$ and by Lemma˜5.1 the graph $G^{\prime}[N[C]]$ contains an induced subgraph $H^{\prime}\in\mathcal{S}\cup\mathcal{T}\cup\mathcal{M}$ with extremities $\{u,v,y\}$. Thus, $H=G^{\prime}[V(H^{\prime})\cup S]$ either
  • belongs to $\mathcal{M}_{k}$ if $H^{\prime}\in\mathcal{M}$,
  • contains an $S_{k}$ as an induced subgraph if $H^{\prime}\in\mathcal{S}$,
  • or contains a $T_{k}$ as an induced subgraph if $H^{\prime}\in\mathcal{T}$.
By applying ˜5.2 and ˜5.3, we get one of the outcomes (i) or (ii) (or we exhibit the wanted induced subgraph $H$) and, as discussed, the number of such steps is at most $n^{2}$. The running time of each individual step is bounded by $\mathcal{O}(m+n)$, as stated in the respective claims. The claimed time complexity follows. ∎
@@ 6 Concluding Remarks and Open Questions
Our main result, Theorem˜1.4, shows that for every integer $k\geqslant 1$, each connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. This raises the question regarding the complexity of computing the path eccentricity of such a graph. While the problem is clearly polynomial for $k=1$, it turns out that the problem is NP-hard for all $k\geqslant 2$. This follows from the fact that the problem of determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96], which are known to be $2P_{2}$-free and, hence, $\{S_{2},T_{2}\}$-free. This, however, says nothing about the complexity of computing a path witnessing the upper bound on path eccentricity guaranteed by Theorem˜1.4.
@@
Is there an algorithmic version of Theorem˜1.4? More precisely: for a fixed integer $k\geqslant 1$ and a connected $\{S_{k},T_{k}\}$-free graph $G$, can we find in polynomial time a path in $G$ with eccentricity less than $k$?
While ˜6.1 is known to have an affirmative answer for $k=1$ (see [She91, BDK00]), it is, to the best of our knowledge, open for all $k\geqslant 2$. Theorem˜1.9 provides a partial answer.
Another result of this paper (Theorem˜1.8) is that every connected $3P_{2}$-free graph $G$ has a longest path that is dominating. This result, along with the observation that there exist graphs that are not $3P_{2}$-free such that every connected induced subgraph has a longest path that is dominating (consider, for instance, paths and cycles), raises the question about a characterization of such graphs.
@@
Is it true that for every integer $k\geqslant 2$, every connected $\{S_{k},T_{k}\}$-free graph has a minimum connected distance-(k1)(k-1) dominating set inducing a $\{S_{1},T_{1}\}$-free graph?